Seven times what number is 1596?
Found 2 solutions by jim_thompson5910, lenny460:
Answer by jim_thompson5910(35256)

"Seven times what number is 1596" --> 7x = 1596

7x = 1596

7x/7 = 1596/7

x = 228

So

"Seven times what number is 1596"

turns into

"Seven times 228 is 1596"


--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by lenny460(1073)

1596 / 7 = 228

The Answer:

228


obuong3@aol.com